OpenStudy (anonymous):

dont need weird mathematics :) \[f(x)=\sqrt{x^2-2x+10}-\sqrt{x^2-2x+2}\] find all \(x \in \mathbb{R}\) such that \(f(x) \in \mathbb{N}\)

OpenStudy (anonymous):

@sauravshakya @ganeshie8 @eliassaab @TuringTest

OpenStudy (nipunmalhotra93):

x=1,\[\frac{ -1-\sqrt{5} }{2 }\], \[\frac{-1+\sqrt{5} }{2 }\]

OpenStudy (nipunmalhotra93):

is that it?

OpenStudy (anonymous):

let me check it :)

OpenStudy (anonymous):

u are close :) but dont tell the answer plz...so others can think on that :)

OpenStudy (nipunmalhotra93):

Ok fine. Thanks :)

OpenStudy (anonymous):

So only when x=1 or are there any other value..... I think yes because irrational-irrational can be integer

OpenStudy (anonymous):

some other values...

OpenStudy (anonymous):

I mean are there any other values of x

OpenStudy (anonymous):

yeah there are.

OpenStudy (anonymous):

dont post ur solution :) just values of x

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So I delete it and send u msg for the vaues of x

OpenStudy (anonymous):

no give the values of x here :)

OpenStudy (anonymous):